# Advancing Cross-domain Discriminability in Continual Learning of Vison-Language Models

### v1

@@ Appendix F Detailed performance of RAIL with order-I
Appendix G Comparison of different methods on few-shot MTIL setting
In this section, we evaluate the performance of RAIL in 5-shot MTIL setting, comparing it against the performance of baselines reported in [14]. In the context of MTIL, where the domain-identity is provided during testing, RAIL is reduced to the structure of multiple domain-specific classifiers trained on each domain separately. The domain-identity guides the test image to the corresponding classifier for within-domain prediction. As shown in 2, RAIL still surpasses all baseline methods in both primal and dual forms. This superior performance can be attributed to the non-linear projection, which significantly enhances the separability of features extracted by CLIP.
Zero-shot 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 65.3
Fine-tune 30.6 93.5 76.8 65.1 91.7 92.9 83.3 96.6 84.9 65.4 71.3 77.5
Transfer
LwF 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
LwF-VR 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
WiSE-FT 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
ZSCL 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
MoE-Adapter 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
Primal-RAIL 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
Dual-RAIL 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
Average
LwF 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
LwF-VR 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
WiSE-FT 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
ZSCL 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
MoE-Adapter 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
Primal-RAIL 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
Dual-RAIL 36.0 94.2 70.9 58.8 70.6 84.3 88.5 70.3 89.7 66.5 65.8 72.3
Last
LwF 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
LwF-VR 22.9 89.9 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
WiSE-FT 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
ZSCL 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
MoE-Adapter 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
Primal-RAIL 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
Dual-RAIL 36.0 94.8 71.5 64.1 79.5 95.3 88.5 89.4 91.5 74.6 71.3 77.9
Appendix H Comparison of different methods on X-TAIL with order II.
@@ Appendix G Comparison of different methods on few-shot MTIL setting
Appendix I Limitation
RAIL exhibits its superior performance and efficiency for transfering pre-trained VLMs to various domains. A limitation here is that the pre-trained VLM remains frozen, with its feature extraction ability unenhanced during the incremental learning process. A promising direction for future work is to adjust the pre-trained encoder according to new data with low computational cost, thus further boosting RAIL’s performance while maintain its efficiency. Beside, extending RAIL to encompass additional downstream tasks of VLMs, such as image segmentation, can broaden its applicability and enhance its utility in more complex visual understanding scenarios.

### v2

@@ Appendix F Detailed performance of RAIL with order-I
Appendix G Comparison of different methods on X-TAIL with order II.
@@
Appendix H Limitation
RAIL exhibits its superior performance and efficiency for transferring pre-trained VLMs to various domains. A limitation in this approach is that the pre-trained VLM remains fixed, preventing any improvement in its feature extraction capability throughout the incremental learning process. A promising direction for future work is to adjust the pre-trained encoder according to new data with low computational cost, thus further boosting RAIL’s performance while maintain its efficiency. Beside, extending RAIL to encompass additional downstream tasks of VLMs, such as image segmentation, can broaden its applicability and enhance its utility in more complex visual understanding scenarios.